# Weak Memory Model Formalisms: Introduction and Survey

### v1

@@ 9.1. Empirically-Established Properties
The influential “herding cats” work of Alglave et al. [36] robustly and rigorously tests weak memory models against litmus tests. This work included testing real hardware across a range of architectures and vendors, and identified a bug in a processor that resulted in a “Programmer Advice Notice” [38], and has been used in the development of Arm processors [27] and the Linux kernel [32].
Based on this work Arm has led the way in formalising aspects of its architecture [27; 28; 123], including publishing the memory model [196, Chapter B2] and the herdtools suite of applications [39; 23]. Ideally other companies and committees would follow suit, publishing, preferably, a formal model, or at least a representative set of litmus tests (small concurrent programs with an associated reachability condition Section 1.2) that cover both allowed (weak) and forbidden behaviours. This would support easier comparison across models and formalisms, and aid cross-compilation.
However, when focussing purely on empirical evaluation, it is unclear how to achieve complete coverage of all behaviours. This is obviously a difficult problem: given say 10 individual instruction types (loads, stores, fence types, memory ordering constraints, branch points, etc.), across just 2 concurrent processes each with 5 lines of code, there are approximately 10 billion possible litmus tests. This number does not include any attempt to generate the expected results. Furthermore, litmus tests typically do not cover loops, or non-trivial conditional statements. A “minimal”, maintained set of tests and expected results could be kept by an architecture vendor; programmers and formalists together could access this database, with both small litmus tests and larger programs explaining weak behaviours that can arise. This would help formalists check that their definitions and representations are correct, and, with the help of a common notation that abstracts from processor-specific syntax (e.g., [25; 26]), would aid comparison and re-use of results across different architectures and languages.
@@ 9.2. Formally-Derived Properties
One theoretical aspect upon which the language-design community has more or less reached a consensus is that “programs that contain no data-race should behave as if under Sequential Consistency” [16], a property known as Data-Race-Free implies Sequential Consistency, or the Data Race Freedom Guarantee (DRF). This relatively early concept [18; 19] provides a base-level property for programmers of sequential, non-interfering code, similar to the principals underlying the circumstances under which localised pipelined reordering can occur (Section 2.1). However, the result does not necessarily help programmers of low-level systems code that does include data races, such as system-wide locks [32; 88].
What would complement such properties and litmus test conformance is a set of more general properties of programs that can be understood locally [46]; for instance, if a fence is inserted between instructions i and j, can standard ordering be assumed? If we let ‘’ represent strict ordering, not subject to memory model effects, then this property of fences can be expressed as follows.
@@
Property (11) would provide the underlying justification for the earlier program-level understanding of property (10), giving assurance that the insertion of a (full) fence allows reasoning about the separated parts of a program separately (i.e., compositionality). Property (12), for some model-specific definition of “independent”, indicates that low-level parallelism is a factor despite the use of sequential composition. Property (13) states that, for a given formalism, instantiating the memory model by the simplest – Sequential Consistency– reduces the program to a standard trace semantics, represented by . Even this relatively simple property – a sanity check on the definitions and structures used in a formal semantics – is rarely formally proved. A general class of properties of programs at this level would make the task of comparing formalisms more straightforward, and provide a relatively stable basis for programmers and verifiers to understand their task. Reduction to standard operators also opens up the possibility for existing tools and techniques to be applied for reasoning, reducing the need for specialised inference systems [26; 181; 93; 111; 110; 295], and expanding the range of properties and domains that can be addressed.
@@ 10. Current and Future Directions: adapting and contributing to software development
We outline some other research topics related to weak memory models, where software developers must face the consequence of weak memory effects. This includes ongoing work into verification methods and tool support, along with model checkers (Sections 7 and 8); in addition, formalisation of weak memory models can and should also impact other aspects of formal concurrency research, such as verified compiler efforts (e.g. [259; 284]) and data race detection [20; 166]. These topics typically require some abstraction from the specific representation of weak memory to facilitate more standard analyses; we give an overview of some of these below where not previously discussed.
@@ 10.1. Properties of Programs
Linearisability is a widely-accepted correctness criterion for concurrent data structures. In contrast to reachability of final states, linearisability describes the continual interaction of processes operating on a shared data structure. Based on its original definition (which assumed Sequential Consistency), Travkin et al. [280; 281] developed techniques to reduce a TSO program into an SC program, enabling one to utilise existing tools such as the SPIN model checker to test linearisability. Other approaches to adapting linearisability for application to weak memory models propose alternative versions of linearisability [107; 106; 269; 268].
@@
Most existing formalisms and tools for weak memory models handle safety properties only. Verifying liveness properties (also called progress properties) is more challenging even in the Sequential Consistency setting, with one reason being the need to address various fairness assumptions. In the weak-memory setting, works that deal with liveness properties only started to emerge recently [180; 7; 8].
@@
These may emerge from microarchitectural features, including through the exploitation of speculative execution and other mechanisms that cause out-of-order execution. The intersection of vulnerabilities and weak memory models is thought to be a potentially important area to investigate [75], and has been addressed by several authors utilising a range of different representations [283; 236; 108; 139; 91; 87; 90]. The addition of security-focussed instructions, such as Capability Hardware Enhanced RISC Instructions (CHERI), also impacts vulnerability research [299], and will require richer pointer-level models to address in the presence of weak memory.
@@ 10.2. Adapting Existing Code
The specification of concurrent libraries with weak memory is a crucial topic, as concurrent libraries are the building-blocks of real-world concurrent programs. Based on formalisms for defining memory models, much recent effort has extended and applied them to specifying and verifying concurrent libraries. These works focus on various memory models, from TSO [67] and Java [118] to fragments of the C memory model [47; 97; 103; 242; 264]. Large-scale migration of code between hardware platforms has been recently addressed by Beck et al. [51], who developed a tool for automatically porting x86 code to platforms with weaker memory models, such as Arm and RISC-V. Such tools utilises fence insertion and model checking techniques, and also employs heuristics to improve scalability.
@@
For TSO and PSO, there are fence insertion algorithms based on modelling the write-buffers [6; 190; 191; 192] and on predicate abstraction techniques [5; 10; 11]. In the case of structured programs, Taheri et al. [276] developed a polynomial-time greedy algorithm for fence insertion, and also demonstrated the NP-hardness of the minimum fence insertion problem with multiple types of fences. Many other approaches to fence insertion have been taken [29; 52; 66; 177; 161; 214; 215], with the most recent being [225; 265]. Finding the minimal set remains an open problem, however, as does finding the most computationally efficient set.
@@ 10.3. Programming Styles and Hardware Variants
Transactional memory [140] defines synchronisation mechanisms that allow the programmer to specify atomic “blocks” of instructions. While these blocks should appear atomic, on the hardware level they can still occur in parallel. When two atomic blocks attempt to change the same variable and a clash arises, the transactional memory system can safely revert to a consistent state. Naturally, transactional memory has interplay with memory consistency models [82; 115]. Most recently, Dalvandi and Dongol [98] devised a semantics of transactional memory under the release-acquire fragment of the C memory model, and developed a program logic based on their semantics.
@@
Persistent memory [131] is a relatively new type of physical memory that maintains its state even after system failures, with the aim of avoiding the loss of data. Updates to memory are held in the volatile memory first, and they are flushed to the persistent memory in some order, during which the persistent memory may temporarily be in some inconsistent states. In the case of a power outage, one needs to ensure that the entire memory system can be reverted back to a consistent state. These issues, related to memory consistency and cache coherence, have been addressed by formalisations of weak-memory models [245; 243; 81; 168; 287; 170; 244].
@@ 11. Conclusions
Often this issue can be addressed by having programming abstractions between the assembler and the high-level code; however this is not possible for programmers at the interface of hardware and software, for instance, operating systems developers [32]. In this context safety, security and performance are all of high importance, and a clear place where a rigorous formal model of the behaviour of multicore systems is invaluable [75]. Since low-level details are almost unavoidable when addressing weak memory model concerns, further refinement of formalisms to include detailed instruction semantics [40] would assist binary-level analysis (e.g., [251; 138]).
@@
Researchers at Arm have adopted the axiomatic style to specify their model, which is made concrete via conformance to a large set of litmus tests. This essentially makes the axiomatic style the de facto standard in describing weak memory models. Typically new semantics styles are validated by showing conformance to the expected outcomes of a set of litmus tests, of which canonical sets exist for both Arm and RISC-V. The weakness of the axiomatic style is that it does not lend itself directly to a straightforward semantics for programs. While memory models can be compared, and general properties of memory models described [36], one cannot easily take an abstract program and apply a set of (denotational) rules to obtain the set of traces for a program written in some abstract syntax. This is addressed by, for instance, the Promising semantics [165], an operational semantics which includes axiomatic-style relations. However this semantics, partly due to capturing a wide-range of weak behaviours, is quite complex to apply to even simple programs. Explanations based on microarchitectural features are usually more straightforward to describe and implement, but less general and less transferrable to the software level.
@@ Acknowledgements.
- Lamport [1979] Lamport. 1979. How to make a multiprocessor computer that correctly executes multiprocess programs. IEEE Trans. Comput. C-28, 9, 690–691.

### v2

@@ 9.1. Empirically-Established Properties
The influential “herding cats” work of Alglave et al. [2014b] robustly and rigorously tests weak memory models against litmus tests. This work included testing real hardware across a range of architectures and vendors, and identified a bug in a processor that resulted in a “Programmer Advice Notice” [Arm Ltd., 2011], and has been used in the development of Arm processors [Alglave et al., 2021] and the Linux kernel [Alglave et al., 2018].
Based on this work Arm has led the way in formalising aspects of its architecture [Alglave et al., 2021, 2024; Fox et al., 2023], including publishing the memory model [Ltd., 2024, Chapter B2] and the herdtools suite of applications [Arm Ltd., 2024; Alglave, 2020]. Ideally other companies and committees would follow suit, publishing, preferably, a formal model, or at least a representative set of litmus tests (small concurrent programs with an associated reachability condition Section 1.2) that cover both allowed (weak) and forbidden behaviours. This would support easier comparison across models and formalisms, and aid cross-compilation.
However, when focussing purely on empirical evaluation, it is unclear how to achieve complete coverage of all behaviours. This is obviously a difficult problem: given say 10 individual instruction types (loads, stores, fence types, memory ordering constraints, branch points, etc.), across just 2 concurrent processes each with 5 lines of code, there are approximately 10 billion possible litmus tests. This number does not include any attempt to generate the expected results. Furthermore, litmus tests typically do not cover loops, or non-trivial conditional statements. A “minimal”, maintained set of tests and expected results could be kept by an architecture vendor; programmers and formalists together could access this database, with both small litmus tests and larger programs explaining weak behaviours that can arise. This would help formalists check that their definitions and representations are correct, and, with the help of a common notation that abstracts from processor-specific syntax (e.g., [Alglave and Cousot, 2016, 2017]), would aid comparison and re-use of results across different architectures and languages.
@@ 9.2. Formally-Derived Properties
One theoretical aspect upon which the language-design community has more or less reached a consensus is that “programs that contain no data-race should behave as if under Sequential Consistency” [Adve and Boehm, 2010], a property known as Data-Race-Free implies Sequential Consistency, or the Data Race Freedom Guarantee (DRF). This relatively early concept [Adve and Hill, 1992, 1993] provides a base-level property for programmers of sequential, non-interfering code, similar to the principals underlying the circumstances under which localised pipelined reordering can occur (Section 2.1). However, the result does not necessarily help programmers of low-level systems code that does include data races, such as system-wide locks [Alglave et al., 2018; Colvin et al., 2024].
What would complement such properties and litmus test conformance is a set of more general properties of programs that can be understood locally [Bargmann and Wehrheim, 2024]; for instance, if a fence is inserted between instructions i and j, can standard ordering be assumed? If we let ‘’ represent strict ordering, not subject to memory model effects, then this property of fences can be expressed as follows.
@@
Property (11) would provide the underlying justification for the earlier program-level understanding of property (10), giving assurance that the insertion of a (full) fence allows reasoning about the separated parts of a program separately (i.e., compositionality). Property (12), for some model-specific definition of “independent”, indicates that low-level parallelism is a factor despite the use of sequential composition. Property (13) states that, for a given formalism, instantiating the memory model by the simplest – Sequential Consistency– reduces the program to a standard trace semantics, represented by . Even this relatively simple property – a sanity check on the definitions and structures used in a formal semantics – is rarely formally proved. A general class of properties of programs at this level would make the task of comparing formalisms more straightforward, and provide a relatively stable basis for programmers and verifiers to understand their task. Reduction to standard operators also opens up the possibility for existing tools and techniques to be applied for reasoning, reducing the need for specialised inference systems [Alglave and Cousot, 2017; Lahav and Vafeiadis, 2015; Coughlin et al., 2021; Doherty et al., 2019, 2022; Wright et al., 2021], and expanding the range of properties and domains that can be addressed.
@@ 10. Current and Future Directions: adapting and contributing to software development
We outline some other research topics related to weak memory models, where software developers must face the consequence of weak memory effects. This includes ongoing work into verification methods and tool support, along with model checkers (Sections 7 and 8); in addition, formalisation of weak memory models can and should also impact other aspects of formal concurrency research, such as verified compiler efforts (e.g. [Ševčík et al., 2013; Trippel et al., 2018]) and data race detection [Adve et al., 1991; Kasikci et al., 2015]. These topics typically require some abstraction from the specific representation of weak memory to facilitate more standard analyses; we give an overview of some of these below where not previously discussed.
@@ 10.1. Properties of Programs
Linearisability is a widely-accepted correctness criterion for concurrent data structures. In contrast to reachability of final states, linearisability describes the continual interaction of processes operating on a shared data structure. Based on its original definition (which assumed Sequential Consistency), Travkin et al. [Travkin et al., 2013; Travkin and Wehrheim, 2014] developed techniques to reduce a TSO program into an SC program, enabling one to utilise existing tools such as the SPIN model checker to test linearisability. Other approaches to adapting linearisability for application to weak memory models propose alternative versions of linearisability [Derrick et al., 2014; Derrick and Smith, 2017; Smith et al., 2020; Smith and Groves, 2019].
@@
Most existing formalisms and tools for weak memory models handle safety properties only. Verifying liveness properties (also called progress properties) is more challenging even in the Sequential Consistency setting, with one reason being the need to address various fairness assumptions. In the weak-memory setting, works that deal with liveness properties only started to emerge recently [Lahav et al., 2021; Abdulla et al., 2023, 2024].
@@
These may emerge from microarchitectural features, including through the exploitation of speculative execution and other mechanisms that cause out-of-order execution. The intersection of vulnerabilities and weak memory models is thought to be a potentially important area to investigate [Cauligi et al., 2022], and has been addressed by several authors utilising a range of different representations [Trippel et al., 2019; Ponce de León and Kinder, 2022; Disselkoen et al., 2019; Griffin and Dongol, 2021; Colvin and Winter, 2020; Colvin, 2023; Colvin and Su, 2025]. The addition of security-focussed instructions, such as Capability Hardware Enhanced RISC Instructions (CHERI), also impacts vulnerability research [Zaliva et al., 2024], and will require richer pointer-level models to address in the presence of weak memory.
@@ 10.2. Adapting Existing Code
The specification of concurrent libraries with weak memory is a crucial topic, as concurrent libraries are the building-blocks of real-world concurrent programs. Based on formalisms for defining memory models, much recent effort has extended and applied them to specifying and verifying concurrent libraries. These works focus on various memory models, from TSO [Burckhardt et al., 2012] and Java [Emmi and Enea, 2019] to fragments of the C memory model [Batty et al., 2013; Dalvandi and Dongol, 2021; Dang et al., 2022; Raad et al., 2019a; Singh and Lahav, 2023]. Large-scale migration of code between hardware platforms has been recently addressed by Beck et al. [2023], who developed a tool for automatically porting x86 code to platforms with weaker memory models, such as Arm and RISC-V. Such tools utilises fence insertion and model checking techniques, and also employs heuristics to improve scalability.
@@
For TSO and PSO, there are fence insertion algorithms based on modelling the write-buffers [Abdulla et al., 2012b; Linden and Wolper, 2010, 2011, 2013] and on predicate abstraction techniques [Abdulla et al., 2012a, 2015b, 2015a]. In the case of structured programs, Taheri et al. [2019] developed a polynomial-time greedy algorithm for fence insertion, and also demonstrated the NP-hardness of the minimum fence insertion problem with multiple types of fences. Many other approaches to fence insertion have been taken [Alglave et al., 2014a; Bender et al., 2015; Burckhardt et al., 2007; Kuperstein et al., 2012; Joshi and Kroening, 2015; Meshman et al., 2014, 2015], with the most recent being [Oberhauser et al., 2021; Singh et al., 2022]. Finding the minimal set remains an open problem, however, as does finding the most computationally efficient set.
@@ 10.3. Programming Styles and Hardware Variants
Transactional memory [Guerraoui and Kapatka, 2010] defines synchronisation mechanisms that allow the programmer to specify atomic “blocks” of instructions. While these blocks should appear atomic, on the hardware level they can still occur in parallel. When two atomic blocks attempt to change the same variable and a clash arises, the transactional memory system can safely revert to a consistent state. Naturally, transactional memory has interplay with memory consistency models [Chong et al., 2018; Dongol et al., 2019]. Most recently, Dalvandi and Dongol [2022] devised a semantics of transactional memory under the release-acquire fragment of the C memory model, and developed a program logic based on their semantics.
@@
Persistent memory [Gogte et al., 2022] is a relatively new type of physical memory that maintains its state even after system failures, with the aim of avoiding the loss of data. Updates to memory are held in the volatile memory first, and they are flushed to the persistent memory in some order, during which the persistent memory may temporarily be in some inconsistent states. In the case of a power outage, one needs to ensure that the entire memory system can be reverted back to a consistent state. These issues, related to memory consistency and cache coherence, have been addressed by formalisations of weak-memory models [Raad et al., 2019b, 2020; Cho et al., 2021; Khyzha and Lahav, 2021; Vafeiadi Bila et al., 2022; Klimis et al., 2024; Raad et al., 2022].
@@ 11. Conclusions
Often this issue can be addressed by having programming abstractions between the assembler and the high-level code; however this is not possible for programmers at the interface of hardware and software, for instance, operating systems developers [Alglave et al., 2018]. In this context safety, security and performance are all of high importance, and a clear place where a rigorous formal model of the behaviour of multicore systems is invaluable [Cauligi et al., 2022]. Since low-level details are almost unavoidable when addressing weak memory model concerns, further refinement of formalisms to include detailed instruction semantics [Armstrong et al., 2019] would assist binary-level analysis (e.g., [Rocha et al., 2022; Gouicem et al., 2022]).
@@
Researchers at Arm have adopted the axiomatic style to specify their model, which is made concrete via conformance to a large set of litmus tests. This essentially makes the axiomatic style the de facto standard in describing weak memory models. Typically new semantics styles are validated by showing conformance to the expected outcomes of a set of litmus tests, of which canonical sets exist for both Arm and RISC-V. The weakness of the axiomatic style is that it does not lend itself directly to a straightforward semantics for programs. While memory models can be compared, and general properties of memory models described [Alglave et al., 2014b], one cannot easily take an abstract program and apply a set of (denotational) rules to obtain the set of traces for a program written in some abstract syntax. This is addressed by, for instance, the Promising semantics [Kang et al., 2017], an operational semantics which includes axiomatic-style relations. However this semantics, partly due to capturing a wide-range of weak behaviours, is quite complex to apply to even simple programs. Explanations based on microarchitectural features are usually more straightforward to describe and implement, but less general and less transferrable to the software level.
@@ Acknowledgements.
- Lamport [1979] L. Lamport. 1979. How to make a multiprocessor computer that correctly executes multiprocess programs. IEEE Trans. Comput. C-28, 9, 690–691.